OpenStudy (anonymous):

One leg of a right triange mesures 15 m and the hypotenuse is 17 m, what is the length of the other leg?

OpenStudy (anonymous):

8

OpenStudy (lstrasmann):

\[a^2 + b^2 = c^2\]\[15^2 + b^2 = 17^2\]\[225 + b^2 = 289\]\[b^2 = 64\]\[b=\sqrt{64}\]\[b=8\]

OpenStudy (anonymous):

you have to use the formula: x^2 + y^2 = r^2. 15^2 + y^2 = 17^2 225 + b^2 = 289 b^2 = 64 b = 64−−√ b = 8

OpenStudy (anonymous):

\[\sqrt{64}\]

OpenStudy (anonymous):

 what is the heigth of the tree